two charges 10 micro columb and 18 micro columb are 3m part in air . find force of replulsion between them ?

Grade:10

1 Answers

Isha chandak
2 years ago
F=kq1q2/r2
q1=10* 10-6 , q2=18* 10-6 ,  r= 3 , k= 9* 109
Putting the above values, we get-
F= 9*10* 10*10-6 * 18*10-6/ 3
F = 54* 10 -2 C